If g1 and g2 denote acceleration due to gravity on the surface of the earth and on a planet respectively, whose mass and radius are twice that of the earth,then
 
1) g1 =2g2
2) g2=2g1
3) g1=g2
4) g1=2g2^2

Grade:9

2 Answers

harshin nechooli
5 years ago
  1. g proportional to mass and inversely proportional to sq of radius of planet so so we can write that if M1 is the mass of first planet that is earth and M2 is the mass of the second planet and R1 steam radius of the first planet that is earth and are you is the radius of the second planet then as the question says
  2. g1÷g2=(M1÷M2)×(R2 sq÷R1 sq) 
  3. that gives g1 = 2×g2
  4.  
Sayantan Garai
5 years ago
g=GM/R²
g α M/R²
g1/g2= M?1R2 ² / M2R1 ²
= M1/M2  R2 ²/ R1 ²
=1/2×4/1
=2
Hence, g1=2g2 (option a)